(D-D'+2)(D+D'-1)=e^x-y -x^2y
1
Expert's answer
2020-11-29T19:15:22-0500

Given differential equation is "(D-D'+2)(D+D-1) = e^{x-y}-x^2y"


C.F. of the equation of type "(D-mD'+k)z = 0" is "u = e^{kx}(y+mx)"


So the C.F. of the equation is, "C.F. = e^{-2x}(y+x) +e^x(y-x)"


P.I. "\frac{}{}""\frac{1}{(D-D'+2)(D+D-1) }( e^{x-y}-x^2y)"

For exponential part,

"\frac{1}{(D-D'+2)(D+D-1) } e^{x-y} = \frac{1}{(1+1+2)(1-1-1) } e^{x-y} = \frac{-1}{4}e^{x-y}"


For polynomial part,

"\frac{1}{(D-D'+2)(D+D-1) }( x^2y) = \frac{1}{D^2-D'^2+2D-2}x^2y"


"= -2(1 - \frac{D^2-D'^2+2D}{2})^{-1} x^2y"


"=-2 (1+(\frac{D^2-D'^2+2D}{2})+(\frac{D^2-D'^2+2D}{2})^2 + .....)x^2y"

Removing the terms which will give me zero, "i.e. D^3, D'^2......."

"= -2(1+(\frac{D^2-D'^2+2D}{2}) +D^2)x^2y"

"= -(x^2y+\frac{1}{2}(2y+4xy)+\frac{1}{4}\times 8 ) = -(4+2y+4xy+2x^2y)"


So, Complete solution is,

"u(z) = e^{-2x}(y+x) +e^x(y-x)-\frac{1}{4}e^{x-y}-(4+2y+4xy+2x^2y)"